# Multiplication of 0-1 matrices via clustering

### v2

@@ 3 An approximate arithmetic matrix product of 0-1 matrices
We slightly extend $APPROXMMCLUS(A,B,\ell)$ in order to obtain a preprocessing for answering queries about single entries of the arithmetic matrix product of $A$ and $B.$
procedure $PREPROCMMCLUS(A,B,\ell)$
Input: Two 0-1 matrices $A$ and $B$ of sizes $p\times q$ and $q\times r,$ respectively, where $p\geq r,$ and a positive integer $\ell$ not exceeding $p$.
Output: The $p\times r$ matrix $D$ returned by $APPROXMMCLUS(A,B,\ell)$, and for $1\leq i\leq p$, the set of coordinate indices $ind(A,i)$ on which $A_{i*}$ differs from its cluster center.
  1. 1.
    Run $APPROXMMCLUS(A,B,\ell)$.
  2. 2.
    For $1\leq i\leq p,$ determine the set $ind(A,i)$ of coordinate indices on which $A_{i*}$ differs from $cen_{\ell}(A_{i*})$.
@@
$PREPROCMMCLUS(A,B,\ell)$ can be implemented in
$O(pq\ell+pr+T(\ell,q,r))$ time.
@@
Recall that $p\geq r.$ Step 1 can be done in $O(pq\ell+pr+T(\ell,q,r))$ time by Lemma 2. Step 2 can be easily implemented in $O(pq)$ time. ∎
@@
procedure $QUERYMMCLUS(A,B,\ell,i,j)$
Input: The preprocessing done by $PREPROCMMCLUS(A,B,\ell)$ for 0-1 matrices $A$ and $B$ of sizes $p\times q$ and $q\times r,$ respectively, where $p\geq r,$ $\ell\in[p]$, and two query indices $i\in[p]$ and $j\in[r].$
Output: The inner product $C_{ij}$ of the $i$-th row $A_{i*}$ of $A$ and the $j$-th column $B_{*j}$ of $B.$
  1. 1.
    Set $C_{ij}$ to the entry $D_{ij}$ of the matrix $D$ computed by
    $APPROXMMCLUS(A,B,\ell)$ in $PREPROCMMCLUS(A,B,\ell,k)$.
  2. 2.
    For $m\in ind(A,i)$ do
    1. (a)
      If the $m$-th coordinate of the center assigned to $A_{i*}$ is 00 and $B_{mj}=1$ then $C_{ij}\leftarrow C_{ij}+1$.
    2. (b)
      If the $m$-th coordinate of the center assigned to $A_{i*}$ is $1$ and $B_{mj}$ is also $1$ then $C_{ij}\leftarrow C_{ij}-1$.
@@
$QUERYMMCLUS(A,B,\ell,i,j)$ is correct, i.e., the final value of $C_{ij}$ is the inner product of the $i$-th row $A_{i*}$ of $A$ and the $j$-th column $B_{*j}$ of $B.$
@@
$C_{ij}$ is initially set to $D_{ij}$, which is the inner product of the center assigned to $A_{i*}$ and $B_{*j}.$ Then, $C_{ij}$ is appropriately corrected by increasing or decreasing with $1$ for each coordinate index $m\in ind(A,i)$ which contributes $1$ to the inner product of $A_{i*}$ and $B_{*j}$ and 00 to the inner product of the center of $A_{i*}$ and $B_{*j}$ or vice versa. ∎
@@
$QUERYMMCLUS(A,B,\ell,k,i,j)$ takes $O(\lambda(A,\ell,row))$ time.
@@
Recall that $2\lambda(A,\ell,row)$ is an upper bound on the maximum Hamming distance between a row of $A$ $B$ and its center in the $\ell$-center clustering computed by $APPROXMMCLUS(A,B,\ell)$ in $PREPROCMMCLUS(A,B,\ell)$. Recall also that $p\geq r.$ Step 1 takes $O(1)$ time. Since the $m$-th coordinate in the centers can be accessed in the matrix $A^{\prime}$ computed by $APPROXMMCLUS(A,B,\ell)$, each of the two substeps in the block of the loop in Step 2 can be done in $O(1)$ time. Finally, since $|ind(A,j)|\leq 2\lambda(A,\ell,row)$, the block is iterated at most $2\lambda(A,\ell,row)$ times. Consequently, the whole Step 2 takes $O(\lambda(A,\ell,row))$ time. ∎
By putting Lemmata 3, 4, and 5 together, and using the straightforward $O(sqt)$-time algorithm to multiply matrices of size $s\times q$ and $q\times t,$ we obtain our next main result for $p\geq r.$ The case $p\leq r$ reduces to the case $p\geq r$ by $(AB)^{\top}=B^{\top}A^{\top}.$ Recall that then the number of rows in $B^{\top}$, which is $r$, is not less than the number of columns in $A^{\top}$, which is $p$. Also, we have $\lambda(B^{\top},k,row)=\lambda(B,k,col).$ We simply run $PREPROCMMCLUS(B^{\top},A^{\top},k)$ and
$QUERYMMCLUS(B^{\top},A^{\top},k,j,i)$ instead.
@@ 4 The exact arithmetic matrix product of 0-1 matrices
Let $A$ and $B$ be two 0-1 matrices of sizes $p\times q$ and $q\times r,$ respectively. Given parameters $\ell\in[p]$ and $k\in[r],$ the arithmetic matrix product of $A$ and $B$ can be computed by a simple deterministic algorithm in $O(pq\ell+pr\lambda(A,\ell,row))$ time if $p\geq r$ or $O(rqk+pr\lambda(B,k,col))$ time if $p\leq r.$
There is however a slightly better way of obtaining a simple deterministic algorithm for the arithmetic matrix product of two 0-1 matrices via $\ell$-center clustering of the rows of the first matrix or $k$-center clustering of the columns of the second matrix. The idea is to use the aforementioned technique of traversing an approximate minimum spanning tree of the rows of the first matrix or the columns of the second matrix in an appropriate Hamming space in order to compute a row or column of the product matrix [4, 11, 15]. The technique easily generalizes to 0-1 rectangular matrices. We shall use the following procedure and lemma in the spirit of [4, 11, 15].
procedure $MMST(A,B,T)$
Input: Two matrices $A$ and $B$ of sizes $p\times q$ and $q\times r$, respectively, and a spanning tree $T$ of the rows of $A$ in the Hamming space $\{0,1\}^{q}.$
Output: The arithmetic matrix product $C$ of $A$ and $B$.
  1. 1.
    Construct a traversal (i.e., a non-necessarily simple path visiting all vertices) $U$ of $T.$
  2. 2.
    For any pair $A_{m*}$, $A_{i*}$, where the latter row follows the former in the traversal $U,$ compute the set $diff(m,i)$ of indices $h\in[q]$ where $A_{ih}\neq A_{mh}$.
  3. 3.
    For $j=1,\ldots,r$, iterate the following steps:
    1. (a)
      Compute $C_{sj}$ where $A_{s*}$ is the row of $A$ from which the traversal $U$ of $T$ starts.
    2. (b)
      While following $U$, iterate the following steps:
      1. i.
        Set $m,\ i$ to the indices of the previously traversed row of $A$ and the currently traversed row of $A$, respectively.
      2. ii.
        Set $C_{ij}$ to $C_{mj}$.
      3. iii.
        For each $h\in diff(m,i)$, if $A_{ih}B_{hj}=1$ then set $C_{ij}$ to $C_{ij}+1$ and if $A_{mh}B_{hj}=1$ then set $C_{ij}$ to $C_{ij}-1$.
@@
First, we shall prove that $MMST(A,B,T_{A})$ computes the arithmetic matrix product of $A$ and $B$ in time $O(pq+qr+r\times\mathrm{ham}(T_{A})).$ The correctness of the procedure $MMST$ follows from the correctness of the updates of $C_{ij}$ in the block of the inner loop, i.e., in Step 3(b). Step 1 of $MMST(A,B,T_{A})$ can be done in $O(p)$ time while Step 2 requires $O(pq)$ time. The first step in the block under the outer loop, i.e., computing $C_{sj}$ in Step 3(a), takes $O(q)$ time. The crucial observation is that the second step in this block, i.e., Step 3(b), requires $O(p+\mathrm{ham}(T_{A}))$ time. Simply, the substeps (i), (ii) take $O(1)$ time while the substep (iii) requires $O(|diff(m,i)|+1)$ time. Since the block is iterated $r$ times, the whole outer loop, i.e., Step 3, requires $O(qr+pr+r\mathrm{ham}(T_{A}))$ time. Thus, $MMST(A,B,T_{A})$ can be implemented in time $O(pq+qr+rp+r\times\mathrm{ham}(T_{A})).$
Similarly, we can run $MMST(B^{\top},A^{\top},T_{B})$ to obtain the transpose of the arithmetic matrix product of $A$ and $B.$ So, to obtain the lemma, we can alternate the steps of $MMST(A,B,T_{A})$ and $MMST(B^{\top},A^{\top},T_{B})$, and stop whenever any of the calls is completed. ∎
@@
We determine an $\ell$-center clustering of the rows of $A$ in $\{0,1\}^{q}$ of maximum cluster radius not exceeding $2\lambda(A,\ell,row)$ in $O(pq\ell)$ time by employing Fact 1. Similarly, we construct a $k$-center clustering of the columns of $B$ in $\{0,1\}^{q}$ of maximum cluster radius not exceeding $2\lambda(B,k,col)$ in $O(rqk)$ time. The centers in both aforementioned clusterings are some rows of $A$ and some columns of $B$, respectively, by the specification of the method in [16]. Hence, the $\ell$-center clustering gives rise to a spanning tree $T_{A}$ of the rows of $A$ with all members of a cluster being pendants of their cluster center and the centers connected by a path of length $\ell-1.$ The Hamming cost of $T_{A}$ is at most $(p-\ell)2\lambda(A,\ell,row)+(\ell-1)q.$ Similarly, we obtain a spanning tree $T_{B}$ of the columns of $B$ having the Hamming cost not exceeding $(r-k)2\lambda(B,k,col)+(k-1)q.$ The theorem follows from Lemma 6 by straightforward calculations. ∎
5 Extensions

### v3

@@ 3 An approximate arithmetic matrix product of 0-1 matrices
The heavy three-linear terms $pq\ell$ and rqkrqk in the upper time bounds in Theorem 3.1 originate from the upper time bound on the center clustering problem in Fact 1 and the time taken by matrix multiplication. We can get rid of these terms by using a faster randomized algorithm for approximate center clustering according to Fact 2, both on the rows of the first matrix $A$ and the columns of the second matrix $B$, in a way similar to [AC11]. Then, only the small matrices induced by the centers of the rows of $A$ and the columns of $B$, respectively, need to be multiplied. In this way, the terms $pq\ell$ and rqkrqk are replaced by the presumably smaller term $\ell qk$ in the upper time bound at the cost of increasing the additive error.
procedure MMCLUSMMCLUS-$R$-$Approx(A,B,\ell,k,\epsilon)$ Input: Two 0-1 matrices $A$ and $B$ of sizes $p\times q$ and $q\times r,$ respectively, positive integers $\ell,\ k$ not exceeding p,r,p,\ r, respectively, and a real number $\epsilon\in(0,1/2).$ Output: A $p\times r$ matrix $D$, where for $1\leq i\leq p$ and $1\leq j\leq r,$ $D_{ij}$ is an approximation of the inner product $C_{ij}$ of the $i$-th row $A_{i*}$ of $A$ and the $j$-th column $B_{*j}$ of $B.$ 1. Determine an approximate $\ell$-center clustering of the rows of the matrix $A$ in $\{0,1\}^{q}$ by using Fact 2. For each row $A_{i*}$ of $A$, set $cen_{\ell}(A_{i*})$ to its closest center, with ties broken arbitrarily. 2. Determine an approximate $k$-center clustering of the columns of the matrix $B$ in $\{0,1\}^{q}$ by using Fact 2. For each column $B_{*j}$ of $B$, set $cen_{k}(B_{*j})$ to its closest center, with ties broken arbitrarily. 3. Form the $\ell\times q$ matrix $A^{\prime},$ where the $i^{\prime}$-th row is the $i^{\prime}$-th center in the approximate $\ell$-center clustering of the rows of $A.$ 4. Form the $q\times k$ matrix $B^{\prime},$ where the $j^{\prime}$-th row is the $j^{\prime}$-th center in the approximate $k$-center clustering of the columns of $B.$ 5. Compute the arithmetic $\ell\times k$ matrix product $C^{\prime\prime}$ of $A^{\prime}$ and $B^{\prime}.$ 6. For $1\leq i\leq p$ and $1\leq j\leq r,$ set $D^{\prime}_{ij}$ to $C^{\prime\prime}_{i^{\prime}j^{\prime}}$, where the $i^{\prime}$-th row $A^{\prime}_{i^{\prime}*}$ of $A^{\prime}$ is $cen_{\ell}(A_{i*})$ and the $j^{\prime}$-th column $B^{\prime}_{*j^{\prime}}$ of $B^{\prime}$ is $cen_{k}(B_{*j})$.
@@
Let $C$ stand for the arithmetic product of $A$ and $B.$ For $1\leq i\leq p$ and $1\leq j\leq r,$ $|C_{ij}-D^{\prime}_{ij}|\leq(2+\epsilon)(\lambda(A,\ell,row)+\lambda(B,k,col))$ holds.
@@
For $1\leq i\leq p$ and $1\leq j\leq r,$ $D^{\prime}_{ij}$ is the inner product of $cen_{\ell}(A_{i*})$, where $\mathrm{ham}(A_{i*},cen_{\ell}(A_{i*}))\leq(2+\epsilon)\lambda(A,\ell,row)$, with $cen_{k}(B_{*j})$, where
$\mathrm{ham}(B_{*j},cen_{k}(B_{*j}))\leq(2+\epsilon)\lambda(B,k,col)$ by Fact 2. Hence, $C_{ij}$, which is the inner product of $A_{i*}$ with $B_{*j},$ can differ at most by $(2+\epsilon)(\lambda(A,\ell,row)+\lambda(B,k,col))$ from $D^{\prime}_{ij}.$
Lemma 4
MMCLUSMMCLUS-$R$-$Approx(A,B,\ell,k,\epsilon)$ can be implemented in time
$O(p\log p(q+\ell)/\epsilon^{2}+r\log r(q+k)/\epsilon^{2}+pr+T(\ell,q,k))$.
Proof
In Step 1, the $\ell$ centers of the rows of $A$ can be found in $O(p\log p(q+\ell)/\epsilon^{2})$ time by Fact 2. Also, the $\ell$ clusters induced by the centers can be formed in $O(p\log p(q+\ell)/\epsilon^{2})$ time in order to approximate minimum-diameter $\ell$-clustering of the rows of $A$; see the proof of Fact 2 in [KLP26]. Importantly, each member of such a cluster is within a Hamming distance not exceeding $(2+\epsilon)\lambda(A,\ell,row)$ from its center. Symmetrically, Step 2 takes $O(r\log r(q+k)/\epsilon^{2})$ time. Steps 3 and 4 can be done $O(\ell q)$ time and $O(kq)$ time, respectively, which is $O(T(\ell,q,k))$ time. Finally, Step 5 takes $T(\ell,q,k)$ time while Step 6 can be done in $O(pr)$ time. Thus, the overall time is $O(p\log p(q+\ell)/\epsilon^{2}+r\log r(q+k)/\epsilon^{2}+pr+T(\ell,q,k))$.
Lemmata 3 and 4 yield our second theorem.
Theorem 3.2
Let $A$ and $B$ be two 0-1 matrices of sizes $p\times q$ and $q\times r,$ respectively. There is a simple randomized algorithm which for any $\epsilon\in(0,1/2)$ provides an approximation of all entries of the arithmetic matrix product of $A$ and $B$ within an additive error of w.h.p. at most $(2+\epsilon)(\lambda(A,\ell,row)+\lambda(B,k,col))$ in time $O(p\log p(q+\ell)/\epsilon^{2}+r\log r(q+k)/\epsilon^{2}+\ell qk+pr)$.
We can apply $MMCLUS$-$Approx(A,B,\ell)$ to obtain a preprocessing for answering queries about single entries of the arithmetic matrix product of $A$ and $B.$
procedure $MMCLUS$-$Preproc(A,B,\ell)$ Input: Two 0-1 matrices $A$ and $B$ of sizes $p\times q$ and $q\times r,$ respectively, where $p\geq r,$ and a positive integer $\ell$ not exceeding $p$. Output: The $p\times r$ matrix $D$ returned by $MMCLUS$-$Approx(A,B,\ell)$, and for $1\leq i\leq p$, the set of coordinate indices $ind(A,i)$ on which $A_{i*}$ differs from the closest center. 1. Run MMCLUSMMCLUS-$Approx(A,B,\ell)$. 2. For $1\leq i\leq p,$ determine the set $ind(A,i)$ of coordinate indices on which $A_{i*}$ differs from $cen_{\ell}(A_{i*})$.
Lemma 5
MMCLUSMMCLUS-$Preproc(A,B,\ell)$ can be implemented in $O(pq\ell+pr+T(\ell,q,r))$ time.
Proof
Recall that $p\geq r.$ Step 1 can be done in $O(pq\ell+pr+T(\ell,q,r))$ time by Lemma 2. Step 2 can easily be implemented in $O(pq)$ time.
@@
procedure $MMCLUS$-$Query(A,B,\ell,i,j)$ Input: The preprocessing done by $MMCLUS$-$Preproc(A,B,\ell)$ for 0-1 matrices $A$ and $B$ of sizes $p\times q$ and $q\times r,$ respectively, where $\geq r, [p]\ell\in[p$], and two query indices $i\in[p]$ and $j\in[r].$ Output: The inner product $C_{ij}$ of the $i$-th row $A_{i*}$ of $A$ and the $j$-th column $B_{*j}$ of $B.$ 1. Set $C_{ij}$ to the entry $D_{ij}$ of the matrix $D$ computed by $MMCLUS$-$Approx(A,B,\ell)$ in $MMCLUS$-$Preproc(A,B,\ell,k)$. 2. For $m\in ind(A,i)$ do (a) If the $m$-th coordinate of the center assigned to $A_{i*}$ is 0 and $B_{mj}=1$ then $C_{ij}\leftarrow C_{ij}+1$. (b) If the $m$-th coordinate of the center assigned to $A_{i*}$ is $1$ and $B_{mj}$ is also $1$ then $C_{ij}\leftarrow C_{ij}-1$.
@@
MMCLUSMMCLUS-$Query(A,B,\ell,i,j)$ is correct, i.e., the final value of $C_{ij}$ is the inner product of the $i$-th row $A_{i*}$ of $A$ and the $j$-th column $B_{*j}$ of $B.$
@@
$C_{ij}$ is initially set to $D_{ij}$, which is the inner product of the center assigned to $A_{i*}$ and $B_{*j}.$ Then, $C_{ij}$ is appropriately corrected by increasing or decreasing with $1$ for each coordinate index $m\in ind(A,i)$ which contributes $1$ to the inner product of $A_{i*}$ and $B_{*j}$ and 0 to the inner product of the center of $A_{i*}$ and $B_{*j}$ or vice versa.
@@
MMCLUSMMCLUS-$Query(A,B,\ell,i,j)$ takes $O(\lambda(A,\ell,row))$ time.
@@
Recall that $2\lambda(A,\ell,row)$ is an upper bound on the maximum Hamming distance between a row of $A$ and the closest center in the $\ell$-center clustering computed by $MMCLUS$-$Approx(A,B,\ell)$ in $MMCLUS$-$Preproc(A,B,\ell)$. Recall also that $p\geq r.$ Step 1 takes $O(1)$ time. Since the $m$-th coordinate in the centers can be accessed in the matrix $A^{\prime}$ computed by $MMCLUS$-$Approx(A,B,\ell)$, each of the two substeps in the block of the loop in Step 2 can be done in $O(1)$ time. Finally, since $|ind(A,j)|\leq 2\lambda(A,\ell,row)$, the block is iterated at most $2\lambda(A,\ell,row)$ times. Consequently, the whole Step 2 takes $O(\lambda(A,\ell,row))$ time.
By putting Lemmata 5, 6, and 7 together, and using the straightforward $O(sqt)$-time algorithm to multiply matrices of size $s\times q$ and $q\times t,$ we obtain our next main result for $p\geq r.$ The case $p\leq r$ reduces to the case $p\geq r$ by $(AB)^{\top}=B^{\top}A^{\top}.$ Recall that then the number of rows in $B^{\top}$, which is $r$, is not less than the number of columns in $A^{\top}$, which is $p$. Also, we have $\lambda(B^{\top},k,row)=\lambda(B,k,col).$ We simply run $MMCLUS$-$Preproc(B^{\top},A^{\top},k)$ and $MMCLUS$-$Query(B^{\top},A^{\top},k,j,i)$ instead.
@@
Analogously, by using Theorem 3.2 with $\epsilon$ set, say, to $\frac{1}{4}$ instead of Theorem 3.1, we obtain the following randomized variant of Theorem 3.3.
Theorem 3.4
Let $A$ and $B$ be two 0-1 matrices of sizes $p\times q$ and $q\times r,$ respectively. Given parameters $\ell\in[p]$ and $k\in[r],$ the matrices can be preprocessed by a simple randomized algorithm in time $O(p\log p(q+\ell)+r\log r(d+k)+\ell qk+pr)$ such that a query asking for the exact value of a single entry $C_{ij}$ of the arithmetic matrix product $C$ of $A$ and $B$ can be answered in $O(\lambda(A,\ell,row)+\lambda(B,k,col))$ time w.h.p.
@@ 4 The exact arithmetic matrix product of 0-1 matrices
Let $A$ and $B$ be two 0-1 matrices of sizes $p\times q$ and $q\times r,$ respectively. Given parameters $\ell\in[p]$ and $k\in[r],$ the arithmetic matrix product of $A$ and $B$ can be computed by a simple deterministic algorithm in time $O(pq\ell+pr\lambda(A,\ell,row))$ if $p\geq r$ or time $O(rqk+pr\lambda(B,k,col))$ if $p\leq r.$
Similarly, we obtain the following corollary from Theorem 3.4.
Corollary 2
[…]
There is however a slightly better way of obtaining a simple deterministic algorithm for the arithmetic matrix product of two 0-1 matrices via $\ell$-center clustering of the rows of the first matrix or $k$-center clustering of the columns of the second matrix than that of Corollary 1. The idea is to use the aforementioned technique of traversing an approximate minimum spanning tree of the rows of the first matrix or the columns of the second matrix in an appropriate Hamming space in order to compute a row or column of the product matrix [BL, FJLL18, GL03]. The technique easily generalizes to 0-1 rectangular matrices. We shall use the following procedure and lemma in the spirit of [BL, FJLL18, GL03].
procedure MMCLUSMMCLUS-$ST(A,B,T)$ Input: Two matrices $A$ and $B$ of sizes $p\times q$ and $q\times r$, respectively, and a spanning tree $T$ of the rows of $A$ in the Hamming space $\{0,1\}^{q}.$ Output: The arithmetic matrix product $C$ of $A$ and $B$. 1. Construct a traversal (i.e., a non-necessarily simple path visiting all vertices) $U$ of $T.$ 2. For any pair $A_{m*}$, $A_{i*}$, where the latter row follows the former in the traversal $U,$ compute the set $diff(m,i)$ of indices $h\in[q]$ where $A_{ih}\neq A_{mh}$. 3. For $j=1,\ldots,r$ do: (a) Compute $C_{sj}$ where $A_{s*}$ is the row of $A$ from which the traversal $U$ of $T$ starts. (b) While following $U$, do: i. Set m,im,\ i to the indices of the previously traversed row of $A$ and the currently traversed row of $A$, respectively. ii. Set $C_{ij}$ to $C_{mj}$. iii. For each $h\in diff(m,i)$, if $A_{ih}B_{hj}=1$ then set $C_{ij}$ to $C_{ij}+1$ and if $A_{mh}B_{hj}=1$ then set $C_{ij}$ to $C_{ij}-1$.
@@
First, we shall prove that $MMCLUS$-$ST(A,B,T_{A})$ computes the arithmetic matrix product of $A$ and $B$ in time $O(pq+qr+r\times\mathrm{ham}(T_{A})).$ The correctness of the procedure follows from the correctness of the updates of $C_{ij}$ in the block of the inner loop, i.e., in Step 3(b). Step 1 of $MMCLUS$-$ST(A,B,T_{A})$ can be done in $O(p)$ time while Step 2 requires $O(pq)$ time. The first step in the block under the outer loop, i.e., computing $C_{sj}$ in Step 3(a), takes $O(q)$ time. The crucial observation is that the second step in this block, i.e., Step 3(b), requires $O(p+\mathrm{ham}(T_{A}))$ time. Simply, the substeps (i), (ii) take $O(1)$ time while the substep (iii) requires $O(|diff(m,i)|+1)$ time. Since the block is iterated $r$ times, the whole outer loop, i.e., Step 3, requires $O(qr+pr+r\mathrm{ham}(T_{A}))$ time. Thus, $MMCLUS$-$ST(A,B,T_{A})$ can be implemented in $O(pq+qr+rp+r\times\mathrm{ham}(T_{A}))$ time.
Similarly, we can run $MMCLUS$-$ST(B^{\top},A^{\top},T_{B})$ to obtain the transpose of the arithmetic matrix product of $A$ and $B.$ So, to obtain the lemma, we can alternate the steps of $MMCLUS$-$ST(A,B,T_{A})$ and $MMCLUS$-$ST(B^{\top},A^{\top},T_{B})$, and stop whenever any of the calls is completed.
@@
We determine an $\ell$-center clustering of the rows of $A$ in $\{0,1\}^{q}$ of maximum cluster radius not exceeding $2\lambda(A,\ell,row)$ in $O(pq\ell)$ time by employing Fact 1. Similarly, we construct a $k$-center clustering of the columns of $B$ in $\{0,1\}^{q}$ of maximum cluster radius not exceeding $2\lambda(B,k,col)$ in $O(rqk)$ time. The centers in both aforementioned clusterings are some rows of $A$ and some columns of $B$, respectively, by the last part of Fact 1. Hence, the $\ell$-center clustering gives rise to a spanning tree $T_{A}$ of the rows of $A$ with all members of a cluster being pendants of their cluster center and the centers connected by a path of length $\ell-1,$ see Fig. 2. The Hamming cost of $T_{A}$ is at most $(p-\ell)2\lambda(A,\ell,row)+(\ell-1)q.$ Similarly, we obtain a spanning tree $T_{B}$ of the columns of $B$ having the Hamming cost not exceeding $(r-k)2\lambda(B,k,col)+(k-1)q.$ The theorem follows from Lemma 8 by straightforward calculations.
We note that in case $\ell=k$ and $\lambda(A,\ell,row)=\lambda(B,col),$ the upper time bounds in Corollary 1 and Theorem 4.1 coincide with that for the Boolean matrix product established in [AC11].
Figure 2: The spanning tree $T_{A}$ of the rows of AA.
5 Potential extensions